What is the outcomes of group therapy?
There are no sure and positive outcomes. Group therapy will help some people and fail with others.
What percentage of the data falls outside 2 standard deviation of the mean?
4.55% falls outside the mean at 2 standard deviation
The probability of drawing three black cards one at a time with replacement from a standard deck of 52 cards is 1/3x1/2x26/52, which is 0.833.
Which set of events is an example of dependent events?
A dependent event is one that relies on something else to happen.
For instance with a mouse trap the trap closing "depends" on the trigger being activated.
The probability of drawing the first ace is 4 in 52. The probability of getting the second ace is 3 in 51. The probability of getting the third ace is 2 in 50. The probability, then, of drawing three aces is (4 in 52) times (3 in 51) times (2 in 50), which is 24 in 132600, or 1 in 5525, or about 0.0001810
What are the odds of rolling a double one in dice throwing?
Possible outcomes with one cube = 6
Possible outcomes with 2 cubes = (6 x 6) = 36
Probability of a double-one (snake eyes) = 1/36 = 2.78% (rounded)
Odds: 35 to 1 against
1/5 or 0.2
What is the probability of rolling an even number on a die?
Total numbers on a cube = 6
Even numbers on a cube = 3
Probability of rolling an even number on one fair cube = 3/6 = 50% .
A coin is tossed 100 times how many times would you expect to get heads?
The probability of a heads is 1/2. The expected value of independent events is the number of runs times the probability of the desired result. So:
100*(1/2) = 50 heads
9
What is the experimental probability that exactly 3 children in a family of 4 children will be boys?
4/16 or 0.2 or 25%
What are the requirements of a probability distribution?
It is always non-negative.
The sum (or integral) over all possible outcomes is 1.
5:16
You roll 2 number cubes each labeled 1 to 6 what is probabilty you will roll a double?
The probability is based on the second cube having 6 sides. No matter what you roll on the first cube, the chance of rolling "doubles" is 1 in 6! Now if you ask, what is the chance in rolling a specific set of doubles, the chance is 1 in 6 times 1 in 6 times, or 1 in 36 ...
What is the probability of getting exactly 2 heads in 3 coin tosses?
With 3 coin tosses, the possible outcomes are:
HHH, HHT, HTH, HTT, THH, THT, TTH, TTT
There are 8 possible outcomes and 3 of them have 2 heads. Thus:
probability = 3/8 (= 0.375)
Two independent outcomes with constant probabilities.
This means that if a newborn female is 53.8 cm in length, then she is longer than 95% of other newborn females.
The mean of a symmetrical distribution is 40What is its median?
for symmetrical distributions your mean equals the median. that is one of the properties of the symmetrical distribution.
What is the probability of getting a sum less than 6 if you roll two number cubes?
Out of a total of 36 combinations (1-1, 1-2, 1-3, ... 2-1, 2-2, ... 6-6), you have the following favorable possibilities:
1-1
1-2
1-3
1-4
2-1
2-2
2-3
3-1
3-2
4-1
The total is 10 combinations out of 36, so the probability is 10/36.